# Learning to grok: Emergence of in-context learning and skill composition in modular arithmetic tasks

### v1

@@ Abstract
Large language models can solve tasks that were not present in the training set. This capability is believed to be due to in-context learning and skill composition. In this work, we study the emergence of in-context learning and skill composition in a collection of modular arithmetic tasks. Specifically, we consider a finite collection of linear modular functions $z=a\,x+b\,y\;\mathrm{mod}\;p$ labeled by the vector $(a,b)\in\mathbb{Z}_{p}^{2}$. We use some of these tasks for pre-training and the rest for out-of-distribution testing. We empirically show that a GPT-style transformer exhibits a transition from in-distribution to out-of-distribution generalization as the number of pre-training tasks increases. We find that the smallest model capable of out-of-distribution generalization requires two transformer blocks, while for deeper models, the out-of-distribution generalization phase is transient, necessitating early stopping. Finally, we perform an interpretability study of the pre-trained models, revealing the highly structured representations in both phases; and discuss the learnt algorithm.
11footnotetext: Condensed Matter Theory Center, University of Maryland, College Park22footnotetext: Department of Physics, University of Maryland, College Park33footnotetext: Meta AI44footnotetext: Corresponding author55footnotetext: These authors contributed equally66footnotetext: Source Code: https://github.com/ablghtianyi/ICL_Modular_Arithmetic
@@ 1 Introduction
Large language models (LLMs) can perform simple tasks that have not been present in the training set. This ability is usually achieved via in-context learning [5]. More importantly, LLMs can perform an even larger variety of very complex tasks upon appropriate prompting or fine-tuning. The latter ability to perform complex tasks is usually attributed to the following mechanism. First, LLMs learn a large variety of simple tasks and, then, how to compose those skills to form very complex skills [3]. Furthermore, LLMs also exhibit “emergent capabilities” – a sudden emergence of a new complex skill as a function of scale (either model parameters, compute or data) [28, 9]. It is plausible that these sudden performance improvements are due to one or both of these mechanisms. For example, LLMs show grokking on algorithmic tasks [23], which results from the model learning very structured representations [18, 11, 19]. Once these representations emerge, the model abruptly learns how to perform the task.
@@
Figure 1: (a) The dataset. The tasks are labeled by vectors $(a,b)\in\mathbb{Z}_{p}^{2}$. Each table contains examples of $ax+by\,\,\textrm{mod}\,\,p$. A fraction $1-\alpha$ of the examples is blacked out. The remaining examples are flattened into a single “document” in the batch. Each document is organized as a collection of triples $(x,y,ax+by)$ for $x,y$ from the training set (i.e. not blacked out in the table). To leverage training compute more effectively we predict not just the last element, but also other elements colored in red. This is similar to the traditional next-token prediction, with the main difference that we predict each third token ($x$ and $y$ do not have any correlation between each other). Every task appears exactly the same number of times in each batch. (b) Phase diagram for a six-layer model. We find four different phases. (1) in-distribution memorization: model does not generalize on $x,y$ it has not seen for the training and test tasks. (2) in-distribution generalization: model generalizes on $x,y$ it has not seen for the training tasks, but not for test tasks. (3) out-of-distribution memorization: model does not generalize on $x,y$ it has not seen for test tasks, but completely generalizes on train tasks. (4) out-of-distribution generalization: model generalizes on all $x,y$ for both train and test tasks. We focus on investigating phase (4) in more detail. (c) In-context sample complexity. Accuracy of the model in phase (4) as a function of the number of few-shot examples. (d) Representations developed by one of the attention heads in the first layer. These are projections of the embedding of a pair of numbers onto the two largest principal components (PCs) of the internal representation formed after passing through the attention layer and projection matrix. (e) The first 3 principal components of the embeddings partition the $log_{27}$-annotated numbers into separate planes of even and odd groups. The number 0 is special, and is sandwiched between the two planes.
@@ 2 Related Works
Brown et al. [5] first demonstrated that large models performance improves substantially when a few examples of the task at hand are provided at inference time, in the prompt. Akyürek et al. [2], Ahn et al. [1], von Oswald et al. [27] showed that the AM implements emergent first-order optimization on an emergent objective function to solve linear regression tasks. Furthermore, [2] showed that larger models learn to perform Bayesian estimation. Garg et al. [10] demonstrated that transformers can learn several simple classes of functions in context. Kirsch et al. [15] presented how task diversity and model size would affect the ICL performance for unseen tasks using a mixture of modified MNIST datasets. Raventos et al. [24] investigated the relation between task diversity and out-of-distribution ICL ability on linear regression tasks. Lin and Lee [16] identified two operating modes in ICL using a mixture of linear regression tasks, where for the first several shots, the model tries to figure out the correct task vector and later uses it to predict the correct results. Boix-Adserà et al. [4] showed theoretically and experimentally that with enough pre-training data, a transformer model can perform abstract reasoning that a MLP cannot do. Guo et al. [13] showed that transformers can use lower layers to memorize and upper layers to perform ICL in a feature regression setting. It was found in [25] that ICL is a transient phase from the optimization point of view: it goes away once the model is over-trained. Hendel et al. [14], Liu et al. [17] showed that language models form in-context vectors, which can be extracted and used to control model predictions.
@@ 3 Preliminaries
We consider modular arithmetic tasks of the form: $z^{t}_{i}=a^{t}x_{i}+b^{t}y_{i}\;\mathrm{mod}\;p$. We will refer to the coefficients $(a^{t},b^{t})\in\mathbb{Z}_{p}^{2}$ as the task vector. The superscript $t\in[1,p^{2}]$ labels the $p^{2}$ possible tasks. We will refer to $(x_{i},y_{i})\in\mathbb{Z}_{p}^{2}$ as the input vector, which is labeled by the subscript $i\in[1,p^{2}]$.
@@
Unless stated explicitly, we will use $p=29$, the number of heads $H=4$, and embedding dimension $d_{\mathrm{embed}}=512$, with $n_{\mathrm{ctx}}=32$ in-context examples. All models are trained with AdamW optimizer with batch size $B=1024$ for $200$k steps. We have also tied the embedding layer of the model with the readout layer.
@@ 4.1 Transition driven by the number of tasks
Further evidence supporting the two-circuit competition can be observed in panel (d) of Figure 3. The loss curves show a “monotonic $\rightarrow$ non-monotonic $\rightarrow$ monotonic" transition as the task diversity increases. With a minimal number of pre-training tasks, the model primarily engages in memorization, resulting in a monotonically increasing o.o.d. loss curve. As the number of pre-training tasks increases, the loss curve exhibits non-monotonic behavior, indicating competition between two distinct neural circuits. This transient nature of o.o.d. generalization for $n_{\mathrm{i.d.}}=2^{8}$ is a peculiar case where memorization circuits are initially suppressed but eventually prevail. With substantial task diversity, the circuits responsible for generalization take over, culminating in a monotonic loss curve. Similar insights can be derived from examining the monotonicity of the accuracy curves in panel (e).
@@ 5 Interpretability
Figure 5: $d=4$ and $d=2$ models’ performance on k-shot inference, on the grid of inputs $(x,y)\in\mathbb{Z}_{p}^{2}$ (task vector = $(6,6)$). row 1: Models’ predictions on o.o.d. task of the type $(x_{1}\;\;y_{1}\;\;{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}z_{1}}\;\;\cdots\;\;x_{k}\;\;y_{k}\;\;{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}z_{k}}\;\;x\;\;y\;\;{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}?})$. row 2: Analytical plots showing all the possible examples that are re-scalings of previous examples. In other words, we mark the points white whenever $(x,y)=c_{i}(x_{i},y_{i})$ for some $i\leq k$. row 3: Subtract row 2 from row 1, by using correct=1 and incorrect=0. The red points correspond to the examples that the model manages to get right, but are not re-scalings of any of the previous examples. Whereas, blue points are the examples the model misses, despite being a re-scaling of a previous example. This row tells us about the model’s ability to combine the re-scaled in-context examples. Note that $d=4$ model readily learns to combine previous examples, while its $d=2$ counterpart struggles due to its limited capacity.
An important question to consider is what algorithm the model implements to achieve o.o.d. generalization. Here we propose such an algorithm and provide empirical evidence supporting our claims. For simplicity, consider the case with two-shot inference:
$\left(\begin{matrix}x_{1}&y_{1}&\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}{z^{t}_{1}}&x_{2}&y_{2}&\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}{z^{t}_{2}}&x&y&\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}{?}\end{matrix}\right)$
The model learns to (i) re-scale the in-context inputs and (ii) perform linear transformations (over $\mathrm{GF}(p)$). For example, in the above case, the model needs to find the two constants $c_{1}$ and $c_{2}$ such that
$c_{1}(x_{1},\;y_{1})+c_{2}(x_{2},\;y_{2})=(x,\;y)\quad\mathrm{mod}\;p.$ (2)
Once the model has figured out the constants, the result can be simply computed with modular addition: $c_{1}z^{t}_{1}+c_{2}z^{t}_{2}\;\mathrm{mod}\;p$ [11, 19, 18, 29]. Notably, with sequences that provide more demonstrations, the model has the flexibility to combine in-context examples to derive the answer [^1] In an autoregressive setting, the actual position that the model needs to output the prediction is shifted by one token, but this does not change our argument.. However, no matter which specific implementation the model chooses, solving linear systems in $\mathrm{GF}(p)$ to figure out those constants is a must. This requires the model to be able to combine in-context examples and perform modular arithmetic over them, which breaks down to the following three essential skills other than copying information across in-context examples:
@@
    Multiplication: A necessary skill to rescale in-context examples like in eq. 2
@@
    Addition: A necessary skill for combining two examples in eq. 2
In Section 5.1, we demonstrate that while models of both depths, ($d=2,4$) learn skills I, II perfectly; the $d=4$ model outperforms its $d=2$ in skill III – namely combining the in-context examples. We attribute this disparity to the limited capacity of $d=2$ models. Then, in Section 5.2, we show that there are special attention heads that implement the skills I and II. We explicitly show that the structure of these heads deteriorate as pre-training task diversity $n_{\mathrm{i.d.}}$ decreases. We leave the discussion of embedding layers (shown in Figure 1(e)) to Appendix E.
5.1 Model Learns to Combine
We find telling signatures of Equation 2 upon analyzing model predictions with varying numbers of examples in context. Consider the 1-shot scenario, wherein only one demonstration is given to the model. This is equivalent to setting $c_{2}=0$ in Equation 2. In this case, one would expect that the model will only solve the task correctly for examples that obey $(x,y)=c_{1}(x_{1},y_{1})$ for some $c_{1}$. Indeed, this is exactly what we observe in fig. 5 – the model correctly predicts all the targets when the inputs $(x,y)$ are re-scalings of the in-context input $(x_{1},y_{1})$.
For few-shot scenarios, the model does two things. It can predict the targets correctly whenever the query input is proportional to any of the inputs of the previous demonstrations (Figure 5 row 1). Additionally and more importantly, it can correctly solve substantially more than just these cases (Figure 5 row 1,3). This results from the model’s ability to combine in-context examples and use them for prediction. Evidence for this can be seen in the proliferation of red points in row 3 of Figure 5 as more examples are given in-context. The $d=2$ model performs inferior to its $d=4$ counterpart in this latter skill, hinting at an imperfect implementation of the algorithm and the partially correct choice of the set of constants $\{c_{1},c_{2},\cdots\}$.
5.2 Attention Heads Implement Essential Skills
@@
Figure 6: Models that generalize o.o.d. (left) exhibit more structured attention maps. Additionally, the top-$2$ principal components of the features from the corresponding heads also show more structured patterns. The features are computed for sequences with an o.o.d. task vector $(a^{t},b^{t})=(6,6)$, loop over $(x_{i},y_{i})$ at a specific shot while the previous parts of the sequence are fixed. We annotate each PCA plot with the $(\log_{27}{x_{i}},\log_{27}{y_{i}})\,\,\textrm{mod}\,\,p$ pairs. (a) The principal components form a circle of circles where the position of the outer circle is controlled by $x_{i}$. This pattern remains the same for different task vectors or the shot choices; (b) Only plotted pairs with even $\log_{27}x_{i}$, with each $\log_{27}x_{i}$ circled with different colors. The PCA pattern forms a similar double circle as those in (a), with the key difference that those circles depend on task vector choices and the shot choices; (c, d) Models without o.o.d. generalization ability. We pick heads from the first block that corresponds to the first column of (a). We see clearly the structure of attention maps and PCA patterns deteriorate as the task diversity decreases.
@@
In the previous section, we found “black-box" evidence suggesting that the model is implementing an algorithm akin to solving equation eq. 2. Now, we turn to “white-box" interpretability, to identify the components within the transformer model that are responsible for the essential skills outlined at the beginning of Section 5. In figure Figure 6(a, b) we analyze the important attention heads in a model that generalizes o.o.d. We compare them with the heads from a model that does not generalize o.o.d in Figure 6(c, d).
@@
In the bottom panel, we perform principal component analysis (PCA) on the outputs of this head. Specifically, we feed a batched k-shot sequence of the form $(x_{1}\;\;y_{1}\;\;{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}z_{1}}\;\;\cdots\;\;x_{k}\;\;y_{k}\;\;{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}z_{k}}\;\;x\;\;y\;\;{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}z})$, where the first $k$ inputs are fixed and the last input $(x,y)$ is scanned over all $p^{2}$ possible pairs. We concatenate the resulting features from $x$ and $y$, resulting in $p^{2}\times 2d_{\mathrm{embed}}$ batch of features – and perform PCA on this matrix. We project all these $2d_{\mathrm{embed}}$ dimensional features onto the first two principal components. Annotating the pairs $(x,y)$ with $(\log_{27}x,\log_{27}y)$ [^2] We use 27 as the base of logarithm, which is a primitive root of $\mathrm{GF}(29)$, we find a “Clock-of-clocks" – where clocks of period 28 are themselves arranged in a bigger clock of period 28. Number 0 is located at the center of the clocks [^3] This is expected since $\log_{27}0$ is mathematically ill-defined and needs special treatment.. We observe similar clock-of-clocks for concatenated features from $(x,z)$ and $(y,z)$ as well. We refer the reader to Appendix E for further details.
@@
In Figure 6(b), we analyze the head from layer 2 that effectively implements Skill II – it does so by building upon skill 1 and leveraging the features discussed above. The upper panel shows that the highly structured attention map that focuses on the current example as well as the examples preceding the current one. This pattern aligns with the step in our proposed algorithm where the model needs to compare $(x,y)$ pairs across different examples in-context; and selectively re-scale some of them for later combination. Additionally, as the number of in-context demonstrations increases, the attention map indicates that the model begins to focus on multiple examples simultaneously. This suggests that the model might be employing a potential weighted average approach over these examples, as discussed earlier.
@@
By conducting a PCA analysis similar to that in Figure 6(a), we also identify clocks when annotating examples in the $(\log_{27}x,\log_{27}y)$ format. Unlike the previously discussed pattern, the specifics of this "clock-of-clocks" arrangement vary depending on the position and the choice of task vector $(a^{t},b^{t})$. This variability suggests that the head in question is utilizing information about the specific task from the context to appropriately rescale the equations.

### v2

@@ Abstract
Large language models can solve tasks that were not present in the training set. This capability is believed to be due to in-context learning and skill composition. In this work, we study the emergence of in-context learning and skill composition in a collection of modular arithmetic tasks. Specifically, we consider a finite collection of linear modular functions $z=a\,x+b\,y\;\mathrm{mod}\;p$ labeled by the vector $(a,b)\in\mathbb{Z}_{p}^{2}$. We use some of these tasks for pre-training and the rest for out-of-distribution testing. We empirically show that a GPT-style transformer exhibits a transition from in-distribution to out-of-distribution generalization as the number of pre-training tasks increases. We find that the smallest model capable of out-of-distribution generalization requires two transformer blocks, while for deeper models, the out-of-distribution generalization phase is transient, necessitating early stopping. Finally, we perform an interpretability study of the pre-trained models, revealing highly structured representations in both attention heads and MLPs; and discuss the learned algorithms. Notably, we find an algorithmic shift in deeper models, as we go from few to many in-context examples.
11footnotetext: Department of Physics, University of Maryland, College Park22footnotetext: Meta AI33footnotetext: Corresponding author44footnotetext: Source Code: https://github.com/ablghtianyi/ICL_Modular_Arithmetic
@@ 1 Introduction
Large language models (LLMs) can perform simple tasks absent from their training data. This ability is usually achieved via in-context learning [5]. More importantly, LLMs can perform an even larger variety of very complex tasks upon appropriate prompting or fine-tuning. The latter ability to perform complex tasks is usually attributed to the following mechanism. First, LLMs learn a large variety of simple tasks and, then, how to compose those skills to form very complex skills [3]. Furthermore, LLMs also exhibit “emergent capabilities” – a sudden emergence of a new complex skill as a function of scale (either model size, compute or data) [29, 9]. It is plausible that these sudden performance improvements are due to one or both of these mechanisms. For example, LLMs show grokking on algorithmic tasks [24], which results from the model learning very structured representations [18, 11, 20]. Once these representations emerge, the model abruptly learns how to perform the task.
@@
Figure 1: (a) The dataset. The tasks are labeled by vectors $(a,b)\in\mathbb{Z}_{p}^{2}$. Each table contains examples of $ax+by\,\,\textrm{mod}\,\,p$. A fraction $1-\alpha$ of the examples is blacked out; while the remaining examples are flattened into a single “document” in the batch. Each document is organized as a collection of triples $(x,y,ax+by)$ for $x,y$ from the training set (i.e. not blacked out in the table). Our training is similar to the traditional next-token prediction (autoregressive); with the main difference that we predict every third token, which are marked in red ($x$ and $y$ are uncorrelated). Every task appears exactly the same number of times in each batch. (b) Phase diagram for a six-layer model. We find four different phases. (1) in-distribution memorization: The model only performs well on tasks $(a,b)$ and examples $(x,y)$ from the training set – it does not generalize on unseen examples or tasks. (2) in-distribution generalization: model generalizes on unseen examples $(x,y)$ but not on unseen tasks $(a,b)$. (3) out-of-distribution memorization: model generalizes on unseen tasks $(a,b)$ but only for examples $(x,y)$ it has seen during training. (4) out-of-distribution generalization: model generalizes on unseen tasks $(a,b)$ for seen as well as unseen examples $(x,y)$. We focus on investigating phase (4) in more detail. (c) In-context sample complexity. Accuracy of the model in phase (4) as a function of the number of few-shot examples. (d) Representations developed by one of the attention heads in the first layer. These are projections of the embedding of a pair of numbers onto the two largest principal components (PCs) of the internal representation formed after passing through the attention layer and projection matrix. (e) First 3 PCs of embeddings separate $log_{27}$-annotated numbers into even/odd planes, with 0 sandwiched between them.
@@ 2 Related Works
Brown et al. [5] first demonstrated that large models performance improves substantially when a few examples of the task at hand are provided at inference time, in the prompt. Akyürek et al. [2], Ahn et al. [1], von Oswald et al. [28] showed that the AM implements emergent first-order optimization on an emergent objective function to solve linear regression tasks. Furthermore, [2] showed that larger models learn to perform Bayesian estimation. Garg et al. [10] demonstrated that transformers can learn several simple classes of functions in context. Kirsch et al. [15] presented how task diversity and model size would affect the ICL performance for unseen tasks using a mixture of modified MNIST datasets. Raventos et al. [25] investigated the relation between task diversity and out-of-distribution ICL ability on linear regression tasks. Lin and Lee [16] identified two operating modes in ICL using a mixture of linear regression tasks, where for the first several shots, the model tries to figure out the correct task vector and later uses it to predict the correct results. Boix-Adserà et al. [4] showed theoretically and experimentally that with enough pre-training data, a transformer model can perform abstract reasoning that a MLP cannot do. Guo et al. [13] showed that transformers can use lower layers to memorize and upper layers to perform ICL in a feature regression setting. It was found in [26] that ICL is a transient phase from the optimization point of view: it goes away once the model is over-trained. Hendel et al. [14], Liu et al. [17] showed that language models form in-context vectors, which can be used to steer model predictions.
@@ 3 Preliminaries
We consider modular arithmetic tasks of the form: $z^{t}_{i}=a^{t}x_{i}+b^{t}y_{i}\;\mathrm{mod}\;p$. We will refer to the coefficients $(a^{t},b^{t})\in\mathbb{Z}_{p}^{2}$ as the task vector. The superscript $t\in\{1,\cdots,p^{2}\}$ labels the $p^{2}$ possible tasks. We will refer to $(x_{i},y_{i})\in\mathbb{Z}_{p}^{2}$ as the input vector, which is labeled by the subscript $i\in\{1,\cdots,p^{2}\}$.
@@
Unless stated explicitly, we will use $p=29$, the number of heads $H=4$, and embedding dimension $d_{\mathrm{embed}}=512$, with $n_{\mathrm{ctx}}=32$ in-context examples. All models are trained with AdamW optimizer [19] with batch size $B=1024$ for $200$k steps. We have also tied the embedding layer of the model with the readout layer.
@@ 4.1 Transition driven by the number of tasks
Further evidence supporting the two-circuit competition can be observed in panel (d) of Figure 3. The loss curves show a “monotonic $\rightarrow$ non-monotonic $\rightarrow$ monotonic” transition as the task diversity increases. With a minimal number of pre-training tasks, the model primarily engages in memorization, resulting in a monotonically increasing o.o.d. loss curve. As the number of pre-training tasks increases, the loss curve exhibits non-monotonic behavior, indicating competition between two distinct neural circuits. This transient nature of o.o.d. generalization for $n_{\mathrm{i.d.}}=2^{8}$ is a peculiar case where memorization circuits are initially suppressed but eventually prevail. With substantial task diversity, the circuits responsible for generalization take over, culminating in a monotonic loss curve. Similar insights can be derived from examining the monotonicity of the accuracy curves in panel (e).
@@ 5 Interpreting the Learned Algorithms
Figure 5: $d=4$ and $d=2$ models’ performance on k-shot inference, on the grid of inputs $(x,y)\in\mathbb{Z}_{p}^{2}$ (task vector = $(6,6)$). row 1: Models’ predictions on o.o.d. task of the type $(x_{1}\;\;y_{1}\;\;{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}z_{1}}\;\;\cdots\;\;x_{k}\;\;y_{k}\;\;{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}z_{k}}\;\;x\;\;y\;\;{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}?})$. row 2: Analytical plots showing predictions solely based on Modular Regression algorithm. row 3: Subtract row 2 from row 1, by using correct=1 and incorrect=0. The red points correspond to the examples where Ratio Matching does not give the correct predictions but the model predicts correctly. The blue points are examples that the model missed despite Ratio Matching being applicable. This row tells us about the model’s ability to implement Modular Regression by combining the in-context examples. Note that $d=4$ model readily learns to combine previous examples, while its $d=2$ counterpart struggles due to its limited capacity.
We now explain the algorithms implemented by the models to achieve o.o.d generalization. We find that the models implement two distinct algorithms depending on the depth and the number of in-context examples during inference.
Ratio Matching: If there exists $i$ s.t. $c_{i}(x_{i},y_{i})=(x,y)\;\mathrm{mod}\;p$ then ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}z}=c_{i}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}z^{t}_{i}}\;\mathrm{mod}\;p$. This algorithm only works when $y/x=y_{i}/x_{i}\;\mathrm{mod}\;p$ holds for at least one of the in-context examples.
Modular Regression: Find $\{c_{1},c_{2},\dots,c_{k}\}$ s.t. $\sum_{i=1}^{k}c_{i}(x_{i},y_{i})=(x,y)\;\mathrm{mod}\;p$. Then the prediction is ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}z}=\sum_{i=1}^{k}c_{i}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}z_{i}}\;\mathrm{mod}\;p$. This can be viewed as discretized circular regression over $\mathrm{GF}(p)$.
We find telling signatures of these algorithms upon analyzing model predictions with varying numbers of in-context examples (Figure 5). With very few in-context examples, the models implement the Ratio Matching algorithm. As a canonical example, consider the 1-shot scenario $(x_{1}\;\;y_{1}\;\;{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}z_{1}}\;\;x\;\;y\;\;{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}?})$. In this case, Ratio Matching will only solve the task for inputs that obey $(x,y)=c_{1}(x_{1},y_{1})\;\mathrm{mod}\;p$ for some $c_{1}$. Indeed, this is exactly what we observe in Figure 5 for both $d=2,4$ models. With many ($\sim 10$) in-context examples, $d=2$ and $d=4$ models implement different algorithms. The $d=4$ model can combine the in-context examples using the Modular Regression algorithm, leading to near-perfect o.o.d. generalization. On the other hand, the $d=2$ model still uses Ratio Matching and shows sub-optimal performance. We ascribe this to the limited capacity of the $d=2$ models. In other words, the $d=4$ models exhibits an algorithmic shift from Ratio Matching to Modular Regression as them number of in-context examples increases. Whereas, the $d=2$ models shows no such shift.
To implement these algorithms, the model needs to perform linear operations over $\mathrm{GF}(p)$, which can be broken down into the following essential skills (in addition to copying information, which is readily implemented by attention heads).
@@
    Multiplication: A necessary skill to rescale in-context examples in both algorithms
@@
    Addition: A necessary skill for combining in-context examples in Modular Regression algorithm
While both $d=2,4$ models learn skills I, II perfectly; the $d=4$ model outperforms its $d=2$ counterpart in skill III – which helps it combining the in-context examples.
In the remainder of this section, we show the special attention heads that implement skill I (Section 5.1) and MLPs that implement skills II, III (Section 5.2). We explicitly show the deterioration in this structures as pre-training task diversity ($n_{i.d.}$) decreases. We also elucidate the algorithmic shift in the $d=4$ models as the number of in-context examples increases.
5.1 Attention Heads Implement Skill I
@@
Figure 6: Models that generalize o.o.d. (left) exhibit more structured attention maps. Additionally, the top-$2$ principal components of the features from the corresponding heads also show more structured patterns. The features are computed for sequences with an o.o.d. task vector $(a^{t},b^{t})=(6,6)$, loop over $(x_{i},y_{i})$ at a specific shot while the previous parts of the sequence are fixed. We annotate each PCA plot with the $(\log_{27}{x_{i}},\log_{27}{y_{i}})\,\,\textrm{mod}\,\,p$ pairs. (a) The principal components form a circle of circles where the position of the outer circle is controlled by $x_{i}$. This pattern remains the same for different task vectors or the shot choices; (b) Only plotted pairs with even $\log_{27}x_{i}$, with each $\log_{27}x_{i}$ circled with different colors. The PCA pattern forms a similar double circle as those in (a), with the key difference that those circles depend on task vector choices and the shot choices; (c, d) Models without o.o.d. generalization ability. We pick heads from the first block that corresponds to the first column of (a). Clearly, the structure of attention maps and PCA patterns deteriorate as the task diversity decreases.
Figure 7: Cosine-similarities (Equation 2) between layer outputs $\bm{h}^{l}$ at token $z$ position ($\cos(\bm{h}^{l}_{z}(x,y),\bm{h}^{l}_{z}(x^{\prime},y^{\prime}))$, first row) and token $y$ position ($\cos(\bm{h}^{l}_{y}(x,y),\bm{h}^{l}_{y}(x^{\prime},y^{\prime}))$, second row) reveal patterns in the models’ internal representations. For clarity, we only show selected $x$ and $x^{\prime}$ values, where $y$ and $y^{\prime}$ values range from 00 to $28$ between each tick. For the $d=4$ model, kaleidoscopic patterns in the third layer indicate the generation of all possible $y_{i}/x_{i}$ features for subsequent computations. The last layer shows an algorithmic shift from Ratio Matching to Modular Regression. The $d=2$ model also shows the kaleidoscopic pattern in the first layer, while the second layer identifies the relevant $y/x$ features from the in-context examples for matching.
@@
So far we have presented “black-box” evidence suggesting that the model is implementing the proposed algorithms (Figure 5). Now, we turn to “white-box” interpretability, to identify the components within the transformer model that are responsible for the essential skills. In Figure 6(a, b), we analyze the important attention heads in the $d=2$ model. Specifically, we compare the emergent structures in models trained with different pre-training task diversities.
@@
In the bottom panel, we perform principal component analysis (PCA) on the outputs of this head. Specifically, we feed a batched k-shot sequence of the form $(x_{1}\;\;y_{1}\;\;{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}z_{1}}\;\;\cdots\;\;x_{k}\;\;y_{k}\;\;{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}z_{k}}\;\;x\;\;y\;\;{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}z})$, where the first $k$ inputs are fixed and the last input $(x,y)$ is scanned over all $p^{2}$ possible pairs. We concatenate the resulting features from $x$ and $y$, resulting in $p^{2}\times 2d_{\mathrm{embed}}$ batch of features – and perform PCA on this matrix. We project all these $2d_{\mathrm{embed}}$ dimensional features onto the first two principal components. Annotating the pairs $(x,y)$ with $(\log_{27}x,\log_{27}y)$ [^1] We use 27 as the base of logarithm, which is a primitive root of $\mathrm{GF}(29)$, we find a “circle-of-circles” – where circles of period 28 are arranged in a bigger circle of period 28. Number 0 is located at the center of the circles [^2] This is expected since $\log_{27}0$ is mathematically ill-defined and needs special treatment.. We observe similar circle-of-circles for concatenated features from $(x,z)$ and $(y,z)$ as well. We refer the reader to Appendix E for further details.
@@
In Figure 6(b), we analyze the head from layer 2 that effectively “copies” in-context examples. The upper panel shows the highly structured attention map that focuses on the current as well as the preceding examples. This pattern aligns with the Ratio Matching algorithm where the model compares $(x,y)$ pairs across different in-context examples.
@@
In the bottom panel, by conducting a PCA analysis, we again identify circles when annotating examples in the $(\log_{27}x,\log_{27}y)$ format. However, unlike the previously discussed pattern, the specifics of this “circle-of-circles” arrangement vary depending on the position and the choice of task vector $(a^{t},b^{t})$. This variability suggests that the head in question possesses information about the specific task from the context.